OpenStudy (anonymous):

-4(3m-3)+(-10+8m)=-4m+2

OpenStudy (anonymous):

Is this right

OpenStudy (anonymous):

True. Use Pemdas. Parentheses, exponents, multiplication, division, addition, subtraction. -4(3m-3)+(-10+8m) -4(3m)= -12m -4(-3)= 12 -12m+12+(-10)+8m Combine like terms. 12+(-10)= 2 -12m+8m= -4m -4m+2